OpenStudy (anonymous):

Olivia, a resident of Metropolis, pays Metropolis an annual tax of $70 plus 1.5% of her annual income. If Olivia paid $1,390 in tax, what was Olivia’s income?

OpenStudy (anonymous):

Know how to set this problem up?

OpenStudy (anonymous):

that is my problem, having a hard time figuring it out

OpenStudy (anonymous):

what do you have so far?

OpenStudy (anonymous):

a = annual income 70 + .015a = 1390 (I turned the 1.5% into a decimal) .015a = 1390 - 70 .015a = 1320 a = 1320/.015 a = 88,000 check... 70 + .015(88,000) = 1390 70 + 1320 = 1390 1390 = 1390 (correct) annual income is $88,000